OpenStudy (anonymous):

Solve the system for x: 6x-2y=3 5x+5y=4

OpenStudy (solomonzelman):

But matrix is the best :)

OpenStudy (wolf1728):

A) 6x -2y = 3 B) 5x + 5y = 4 Multiply Equation B by -1.2 B) -6x -6y = -4.8 then add Equation A) A) 6x -2y = 3 -8y = -1.8 y = .225 A) 6x -2*.225 = 3 6x -.45 = 3 6x = 3.45 x = .575